OpenStudy (moonlitfate):

Find the equation of the tangent line the the graph of f at the given point: f(v) = (v-4)(v^2-5), at point (-3,-28).

OpenStudy (anonymous):

so get the derivative

OpenStudy (moonlitfate):

@julian25 -- I've done that all ready. \[f'(v) = [1*(v^2-5)]+[(v-4)*2v]\]\[= v^2-5+2v^2-8v\]\[=3v^2-8v-5\]

OpenStudy (anonymous):

once you have the derivative plug in the values given at the beginning of the problem

OpenStudy (anonymous):

Does that make sense? Let me know.

OpenStudy (moonlitfate):

Yes, it makes sense; so, I would do the following: \[f'(28)=3(-3)^2-8(-3)-5\]\[=-8\] Or did I make a mistake?

OpenStudy (anonymous):

I believe that is correct...I don't recall the exact formula but try that answer